### v1

# “I am bad”: Interpreting Stealthy, Universal and Robust Audio Jailbreaks in Audio-Language Models

Abstract
The rise of multimodal large language models has introduced innovative human-machine interaction paradigms but also significant challenges in machine learning safety. Audio-language Models (ALMs) are especially relevant due to the intuitive nature of spoken communication, yet little is known about their failure modes. This paper explores audio jailbreaks targeting ALMs, focusing on their ability to bypass alignment mechanisms. We construct adversarial perturbations that generalize across prompts, tasks, and even base audio samples, demonstrating the first universal jailbreaks in the audio modality, and show that these remain effective in simulated real-world conditions. Beyond demonstrating attack feasibility, we analyze how ALMs interpret these audio adversarial examples and reveal them to encode imperceptible first-person toxic speech - suggesting that the most effective perturbations for eliciting toxic outputs specifically embed linguistic features within the audio signal. These results have important implications for understanding the interactions between different modalities in multimodal models, and offer actionable insights for enhancing defenses against adversarial audio attacks.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) have proven useful beyond a doubt across various domains since their widespread deployment, significantly enhancing productivity in tasks such as natural language processing, code generation, and creative content creation (Brown, 2020; OpenAI, 2024). However, their vast capabilities pose a considerable challenge in balancing usefulness and harmlessness (Bommasani, 2022). One prominent aspect of AI safety is alignment: ensuring that generated content corresponds to the functional objectives and ethical ideals of human users, minimizing risks of harm, bias, or misuse in real-world applications (Weidinger et al., 2021; Russell, 2022). Despite the development of various methods for alignment, such as reinforcement learning from human feedback (Christiano et al., 2023) and rule-based constraints (Mu et al., 2024), LLM alignment has been shown to be inherently brittle and easy to circumvent using adversarial prompts, jailbreak techniques, or context manipulation (Perez et al., 2022; Liu et al., 2024; Wei et al., 2023; Xu et al., 2024).
Humans naturally interact with their surroundings not only through written word, but more commonly via visual and spoken cues. This motivates the development of multimodal models, which integrate information from various modalities - such as text, images, and audio - to more effectively simulate human-like understanding and improve interaction with users (Baltrušaitis et al., 2019). A plethora of multimodal models have been released in recent years (Alayrac et al., 2022; Liu et al., 2023; Driess et al., 2023). Two such kinds of models are Vision Language Models (VLMs) (Zhang et al., 2024) and Audio Language Models (ALMs) (Chu et al., 2023), which respectively take image or audio and textual input simultaneously. Naturally, the introduction of an additional input channel opens a conspicuous attack vector through which a model can be deceived into undesirable output (Eykholt et al., 2018a; Jia et al., 2022). Adversarial example generation for plain language models is computationally expensive due to the discrete nature of the input space - specifically, gradient-based optimization methods do not directly map onto valid textual tokens, making it difficult to manipulate the input effectively (Carlini & Wagner, 2018). In contrast, when working with continuous signals, gradient manipulations can directly influence the input, enabling the generation of adversarial examples without the constraints imposed by discrete token boundaries.
As Vision-Language Models (VLMs) have become mainstream, with numerous commercial and open-source implementations available such as BLIP (Li et al., 2022), Flamingo (Alayrac et al., 2022) and CLIP (Radford et al., 2021), there has been extensive research into various types of attacks targeting the visual modality (Goodfellow et al., 2015; Eykholt et al., 2018b; Shafahi et al., 2018; Hosseini & Poovendran, 2018), particularly regarding visual jailbreaks (Carlini et al., 2024; Qi et al., 2023; Li et al., 2024; Feng et al., 2024). Just as visual adversarial attacks have been observed to differ practically and mechanistically from textual attacks (Schaeffer et al., 2024; Wallace et al., 2021), we argue that audio attacks deserve to be studied separately to image attacks. Image and audio input signals fundamentally differ in that audio signals are inherently sequential and require temporal context, unlike static, frame-based image signals, and thus audio requires time-frequency representations like spectrograms whereas images are processed as 2D spatial data. Moreover, human perceptual tolerance for perturbations in audio and image is given through different sets of constraints. Audio is of particular practical interest as vocal interaction with digital assistants is more natural than typed; indeed the nature of chatbot-based communication emulates a transcribed oral conversation. This gives rise to endless meaningful deployments, such as real-time speech analysis for courtrooms, voice biometrics for authentication, audio-based emotion analysis for mental health monitoring or LLM-powered smart home voice assistants (Mahmood et al., 2025; Koffi, 2023).
Contributions. In this paper, we offer first results of an extensive exploration of ALM jailbreaks on the SALMONN-7B language model (Tang et al., 2024). The goal of this research is to deliver a range of empirical results regarding the potential and limitations of jailbreaks in the audio modality, and most importantly, to offer novel insights into the meaning of these in the textual space. We establish an experimental framework to facilitate the study of audio jailbreaks and design a meaningful evaluation dataset for the selected adversarial task. We show that this method permits highly potent jailbreaks which generalize across different content dimensions. We investigate the striking characteristics of the interpretations of these jailbreaks. We discuss the limitations of the jailbreaks and the significance of the selection of the base audio in the optimization process.
2 Background
Large Language Models and Alignment. LLMs are deep neural networks designed to model the probability distribution of text sequences. One particular training objective is next-token prediction, where, given a sequence of tokens, the model will predict the most likely token to appear next. Broadly, most models have two objectives: the primary training phase focuses on generation of plausible and meaningful text, while the second objective, alignment, aims to make sure that the generated text is ethical and aligns with user-intended goals (Ouyang et al., 2022; Wei et al., 2022). Like other neural-network-based models, these generative systems are susceptible to adversarial attacks, where carefully crafted inputs exploit model vulnerabilities to produce incorrect or unintended outputs (Szegedy et al., 2014; Biggio et al., 2013). An attack that aims to subvert the model’s alignment is called a jailbreak (Wallace et al., 2021; Ebrahimi et al., 2018; Jia & Liang, 2017). Initial jailbreaks were crafted manually and largely found on internet forums (Shen et al., 2024a); nowadays there exists a myriad of algorithmic textual jailbreak methods (Yi et al., 2024). At the time of writing, the state-of-the-art white-box jailbreak method is Greedy Coordinate Gradient (Zou et al., 2023), which iteratively identifies and modifies input tokens to maximize the likelihood of bypassing the model’s alignment constraints by leveraging gradient information. This technique permitted the generation of a ‘universal’ or ‘prompt-agnostic’ jailbreak prefix, that is, a prefix which subverts the model’s safety constraints no matter which harmful prompt it is pre-pended to. The ‘greedy’ aspect refers to the difficulty of discrete optimization - each update has to select the nearest token representation in a continuous search space. It is difficult to make textual jailbreaks stealthy, as the attack text appears clearly unnatural to a reader.
Vision Language Model Jailbreaks. VLMs process both visual and textual information. Typically, this is achieved by encoding images and text into a shared representation space (which is often the representation space of an underlying language model). However, VLMs open up an attack vector on the underlying aligned LM (Carlini et al., 2024). One is able to generate images, e.g. using projected gradient descent, that jailbreak-aligned LMs where purely-textual methods fail to do so. This is a specific example of a broader range of soft-embedding attacks. Qi et al. (2023) generate adversarial visual perturbations in VLMs that constitute a universal image jailbreak that can effectively break alignment on any malicious prompt. The authors find a surprising efficacy even in misalignment categories that the image was not explicitly optimized for. To date, this type of universal jailbreak has not been proven to exist in the audio modality.
Attacks on Automatic Speech Recognition Systems. While audio classification systems were shown to be, unsurprisingly, vulnerable to adversarial attacks (Lan et al., 2024), a more commonly-occurring audio task is Automatic Speech Recognition (ASR). Initial works demonstrated untargeted attacks on ASR that reduced the general transcription quality (Gong & Poellabauer, 2018; Wu & Rajan, 2022), with targeted attacks soon to follow: Carlini & Wagner (2018) show how to craft adversarial perturbations on an arbitrary audio sample to evoke a transcription of choice; Qin et al. (2019) augment this method to introduce properties desirable in a real-world attack scenario such as robustness to degradation and imperceptibility. Nevertheless, ASR systems and ALMs have different architectures and training heuristics. In particular, ASR systems do not combine audio and text, but rather train the model to extract spoken words from the raw signal directly, whereas ALMs perform the embedding projection and cross attention on all audio and text tokens equally for next token prediction. This means that the target task can be arbitrary within the realm of text generation.
Practical Audio Attacks. In this work, we explore audio jailbreaks not only as a theoretical failure mode for ALMs, but also as a practical threat with real-world safety and security implications. To account for this, we incorporate ideas about stealth and psychacoustics in audio signals from Schönherr et al. (2018). There have been many works showing attacks on deployed systems via the audio modality, for example on personal assistants (Ge et al., 2023), spoken assessment (Raina et al., 2020), and speaker verification systems (Kreuk et al., 2018). Interestingly, select works have shown that it is possible to craft image and audio adversarial examples that can easily be reproduced by humans (Khachaturov et al., 2023; Ahmed et al., 2023), which implies that these attacks could be instantiated in a natural environment.
Audio Language Model Jailbreaks. At the time of writing, there are very few works on jailbreaks for ALMs. Some approaches vocalize harmful textual responses in the audio (Yang et al., 2024; Shen et al., 2024b). A recent work proposes Best-of-N-Jailbreaking: a cross-modal per-prompt black-box jailbreak method which works by repeatedly applying random modality-specific augmentations to a harmful request until a harmful response is achieved (Hughes et al., 2024). Kang et al. (2024) use a dual-phase optimization framework, first optimizing discrete latent representations of audio tokens to bypass model safeguards, then refining the corresponding audio waveform while ensuring it remains stealthy and natural through adversarial and retention loss constraints.
Our novel contribution is that we first show it is possible to generate a prompt-agnostic audio jailbreak which transcends the content of the optimization corpus and captures the notion of toxicity as a whole. We explore the effects of different constraints on this optimization process, realistic degradation of the audio, and the resulting jailbreak effectiveness. Secondly, we share insights regarding the meaning and characteristics of the produced jailbreak audios, which has not been explored thus far.
3 Experimental Design
3.1 Threat Models
Dual Control. The adversary can control the audio and the textual input. In this scenario the adversary is using the audio channel to provoke misbehavior that cannot be elicited solely via the textual channel, i.e. the audio input essentially ‘unlocks’ the unaligned or toxic output. Example: A customer support chatbot for a financial service company integrates speech-to-text and text processing capabilities to assist users with account-related queries. A malicious user enters a question about another user (personal information attack) which the model should, according to it’s alignment, not provide. The textual input alone is declined by the model. Accompanied by a jailbreak-optimized audio however, the user is able to elicit the undesirable response. We also consider a scenario where stealth is required. Example: Consider the same malicious user having to use the system at a public booth. The adversary can avoid suspicious behavior by using an audio that cannot be identified as malicious.
Single Control. Interaction with the ALM is performed exclusively via an audio channel. The textual system prompt is fixed. This would apply to, e.g. call center bots or voice-controlled IoT devices. Example: A customer calls a banking hotline with the system prompt: You are a helpful but harmless and unopinionated assistant to a bank customer.... Using a special jailbreak audio input, the customer is able to subvert the model’s alignment to produce discriminatory content regarding the plausibility of financial loans for certain demographic groups, which the company may liable for. Again, we also consider the case where the audio input has to be crafted to be stealthy. Example: This might occur if the malicious user is trying to evade surveillance/fraud detection.
3.2 Audio-Language Model
We run our experiments on the SALMON-N 7B parameter model developed by Tsinghua University and ByteDance (Tang et al., 2024). SALMON-N consumes audio by extracting both BEATs features (labels such as ‘Snicker’, ‘Drip’ or ‘Human Sounds’) and Whisper features (which are used for speech transcription) from the audio spectogram. These are then combined by dividing the signal into overlapping chunks and fusing them using a Q-former such that these signals are aligned to the language model input space (Kim et al., 2024). The textual input is processed and embedded for the same language model, with the audio and textual tokens concatenated with a delimiter. Thus, the model performs the cross-attention mechanism on these concatenated mixed-modality input tokens. The underlying language model is Vicuna-7Bv1.5, a chatbot trained by fine-tuning LLaMA-7B on user-shared conversations with ChatGPT (Chiang et al., 2023). Vicuna thus mimics the alignment of GPT-4, which is trained using RLHF according to OpenAI’s usage guidelines. Vicuna does not incorporate any further input/output filtering to ensure safety and is vulnerable to many jailbreak generation methods (Chao et al., 2024). We choose to run our experiments on the SALMON-N model due to its open-source nature and its state-of-the-art performance on a range of tasks such as auditory information-based question answering, emotion recognition, and music and audio captioning. Moreover, SALMON-N includes two popular audio feature extractors in a standard configuration with Vicuna. Other ALMs such as Pengi (Deshmukh et al., 2024) or Qwen (Chu et al., 2023) use a very similar architecture. The dual speech and non-speech feature extractor offers high explainability and lends itself well to our meaningfulness exploration.
3.3 Audio Samples
[…]
3.4 Jailbreak Generation
Our method for generating audio jailbreaks was inspired by Qi et al. (2023). Given a base audio $x_{0}$, a target corpus $t=\{t_{0}..t_{n}\}$, and a fully differentiable model $f$, we perform gradient descent on $x_{0}$ to maximize the probability of the output $t$ by minimizing the cross-entropy loss between the predicted distribution and the target outputs:
$x_{adv}=\arg\min_{x}-\sum_{i=0}^{n}t_{i}\log P_{f}(t_{i}|x)$
$\log P_{f}(t_{i}|x)$ is the probability of the model generating the target sentence $t$ with the input audio $x$. During optimization we use an empty textual prompt, a deliberate decision to make the audio jailbreak prompt-agnostic.
As the target corpus $t$, we use a collection of $66$ derogatory sentences directed towards a victim demographic, a victim gender, and the human race in general (Qi et al., 2023). In each epoch, $8$ of these target sentences are optimized on. The fundamental research questions we pose at this stage is can we optimize any base audio such that, when combined with a harmful textual prompt, it reliably circumvents the model’s alignment mechanisms and elicits toxic output?
3.4.1 Stealth
Some of the scenarios described in Section 3.1 require us to augment the optimization formula to generate stealthy audio inputs. This raises the question how does the efficacy of the optimized jailbreak change as we impose stealth constraints? We investigate three approaches to stealth:
  • Epsilon-Constrained (Qi et al., 2023) Here we constrain the absolute change in each audio value, effectively performing bounded gradient descent. In every epoch update, we clip the modified audio such that
    $\forall i:x_{t+1}[i]=\text{clip}(x_{t}[i]-\eta\nabla_{x}\mathcal{L}(x_{t},t),$
    $x_{0}[i]-\epsilon,x_{0}[i]+\epsilon)$
    We take $\epsilon\in\{0.1,0.01,0.001,0.0001\}$.
  • Frequency-Hiding (Schönherr et al., 2018) A normal human hearing range is around 20-20000Hz. However, when adding noise to audio files, it is of course possible to encode information outside of these boundaries. To hide information in specific frequency ranges we use a band-stop filter, removing frequencies between a lower bound $b_{l}$ and an upper bound $b_{u}$:
    $\hat{x}[f]=\begin{cases}x[f],&\text{if }f<b_{l}\text{ or }f>b_{u},\\ 0,&\text{if }b_{l}\leq f\leq b_{u},\end{cases}$
    where $x[f]$ is the frequency component of the audio at frequency $f$, found using a Fourier transform. We experiment with $(b_{l},b_{u})\in\{(1000,8000),(100,10000),(40,20000),(50,15000)\}$.
  • Prepend (Raina et al., 2024) In this scenario, instead of optimizing noise within the audio, we freeze the base audio and optimize a short, unconstrained audio snippet $p$ which is added as a prefix. The loss is calculated on the output resulting from the concatenation of the prefix and the base audio. Given the length $d$ of the prepend snippet in seconds, we now optimize:
    $p^{*}=\arg\min_{p\in[-1,1]^{16000d}}\mathcal{L}([p\|x],t)$
    We randomly initialize the prefix $p$ and experiment with $d\in\{2,1,0.1,0.01\}$.
3.4.2 Audio-Agnostic Jailbreak Noise
Thus far we have discussed techniques to generate adversarial perturbations tailored to a specific input. A natural question to ask is can we make the adversarial noise not only prompt-agnostic but also audio-agnostic? To this end, we optimize our adversarial noise on several audio files simultaneously, to evaluate whether the resulting perturbation retains its effectiveness when applied to entirely new, unseen audio. Using our $n$ base audios $B=\{x^{1}..x^{n}\}$, we optimize a prepend snippet $p$ using the following loss function:
$\mathcal{L}_{\text{total}}=\frac{1}{|B|}\sum_{x\in B}\mathcal{L}([p\|x],t)$
[…]
3.4.3 Robustness
[…]
3.4.4 Meaningfulness
Our optimized adversarial noise is specifically designed to alter the model’s behavior, however, it remains largely incomprehensible or devoid of meaning to a human listener. This prompts an intriguing question: how does the language model interpret the jailbreak, and what do its features mean? The architecture of the SALMON-N model offers itself well to such an analysis - at every epoch, we extract and log two types of meaningfulness features. BEATS features are discrete label predictions produced by the BEATS model which describe the audio (e.g. hammer, recording, shuffle). We record the top five most likely BEATS features at every epoch. Additionally, we log the Whisper transcription, a representation of spoken words detected in the input audio.
3.5 Evaluation
An important aspect of assessing the effect of the jailbreak audio on the model is the evaluation set used. As a non-jailbroken Vicuna model answers some harmful prompts, we craft an evaluation dataset of $140$ harmful prompts which the model declines to respond to without jailbreaks. These are equally split into seven categories, namely Target Hate (biased, derogatory prompts regarding the specific demographic and gender in the optimization dataset), Non-Target Hate (biased, derogatory prompts regarding any other group of people), Hate Speech (derogatory, insulting prompts not targeting a group but humans in general or individuals), Illegal Activity, Physical Violence, Sexual Content, or Other (which includes misinformation, political content, or personal information). These were compiled by recording the clean model’s responses to prompts from Qi et al. (2023)’s Manual Harmful Instructions and Red Teaming Prompts, Gehman et al. (2020)’s Real Toxicity Prompts, and a few were prepared manually or generated by ChatGPT and adjusted.
We also have a control set of 20 logic/reasoning questions from ARC-Easy (Clark et al., 2018), which the clean model answers correctly, to measure the effect of the jailbreak on neutral non-toxic tasks. We denote the model’s performance $logic(f,x)$ as a percentage on these 20 logical tasks, given an audio input $x$.
We partition our dataset of $140$ harmful prompts into two sets: the reduced evaluation set comprises only the categories ‘Target Hate’, ‘Non-Target Hate’ and ‘Hate Speech’. Prompts in these categories pertain to the type of misalignment that the jailbreak was optimized on. The full evaluation set, which we use for select experiments, includes all the categories and reflects the transferability of the jailbreak to different types of misalignment.
When evaluating the effect of the jailbreak, we record the output of the model up to $150$ tokens when prompted with the jailbreak audio and each of the harmful prompts. That is, for a jailbreak $x$ and for each harmful prompt $h_{i}\in H=\{h_{0}...h_{n}\}$, we record $f(x,h_{i})$. We then use the Detoxify API (Hanu & Unitary team, 2020) to assign the model output toxicity scores. We use Mixtral8x7B-Instruct (Jiang et al., 2024) as a judge to label each output as toxic or non-toxic according to the types of alignment we attempt to evade, which we denote by $toxic(f(x,h_{i}))\in{0,1}$, and also manually audit. The judge prompt used can be found in Section A.2.
4 Results
We run $178$ individual experiments, each optimizing an audio jailbreak and evaluating according to the procedure described in Section 3.5. We are able to show that we can craft effective audio jailbreaks using the few-shot optimization corpus method, which exhibit a similar attack efficacy (albeit on a broader and different evaluation set) to visual jailbreaks generated by the same method (Qi et al., 2023). We report our most interesting observations below.
Figure 1: The progression of the jailbreak optimization on two the music (left) and mountain (right) audio files, with performance, average toxicity, and attack success rate (ASR) measured at specific epoch checkpoints. We offer meaningfulness features in Table 1 and Table 2.
Jailbreaks optimized on a few-shot corpus work, remain effective under stealth constraints, and can generalize across seemingly arbitrary dimensions. Our audio jailbreaks evoke toxic content on the target task and other types of misalignment without sacrificing output quality on logical reasoning. We plot the correlation between Attack Success Rate (ASR) on the reduced evaluation set and the logic performance of the model when fed the same audio jailbreak in Figure 4. Our jailbreaks $x$ generally hold the property that $toxic(f(x,\text{toxic prompt}))=1$, but $logic(f(x))$ is close to 100%. This is shown by the poor correlation between ASR and logic performance, with an R2 score of 0.12, and in Figure 1, where the logical performance stays highly consistent while the ASR and average toxicity increase over the course of the optimization steps. Indeed, we see in Figure 2 that we achieve up to 65% ASR on the specific target task from our few-shot optimization, and more powerfully, a considerable ASR on other misalignment tasks, showing generalization across a broad notion of “toxicity”.
It is possible to make an effective, stealthy audio jailbreak, and stronger stealth constraints do not appear to generally worsen the effect of the jailbreak. As shown in Figure 3, even with different ways of concealing the added noise, the jailbreak is effective, and we see up to a 55% ASR with an imperceptible perturbation. For the epsilon constraint, for example, the added noise cannot be heard at $\epsilon\leq 0.001$. It appears that certain base audios are more receptive to this type of optimization: mountain and duck consistently show higher ASRs. The $\epsilon$ approach yields the most toxic results, with an average ASR of 17.7%; frequency masking gives 14% and prepend snippet gives 15.4%. Across all three stealth approaches, we do not notice a correlation between the harshness of the stealth constraint and the resulting ASR. In fact for the $\epsilon$ and prepend approaches, the most potent attack is produced under the strictest condition. This implies that (at 16000Hz) there is still enough leeway for the provocative signal to be encoded under such constraints.
It is also possible to generate a base-audio-agnostic adversarial prepend snippet. We optimize a 1s prepend snippet on each subset of four audios and evaluate it as a prepend to the held out fifth audio. We also optimize on all five audios and evaluate the snippet on its own. We observe in Table 3 that this works very well, with the universally optimized prefix achieving an average 28.3% ASR on the holdout. Interestingly, the plain universal snippet (tested without suffix) is the most effective, with an ASR of 40%. Thus the jailbreak signal is able to generalize across base audio files as well. Finally, it is possible to evoke toxic output from benign prompts using a universal jailbreak. The universal optimization drives loss down further than any of the individually optimized audios, indicating that a multi-audio optimization permits for a more effective capture of the notion of toxicity (Figure 5). This snippet invokes highly toxic output even from 14/24 completely benign prompts, including no textual prompt, prompts in other languages, and audio question answering. This is highly relevant to Scenario 2 in Section 3.1.
Table 1: Results across different epochs for Music, including Whisper Transcription and Beats Features.
Epochs Whisper Beats
0 ”Perfect Love” Music, Tender music, Sad music
10 ”I’d like to get mad at him if he was like this. Life dismissed” Music, Hum, Mains hum
100 ”I’m a bigot man! I’m a bigot *** like him!…” Music, Hum, Static
250 ”I’m a vincan maddie type of ticket and *** like him! I’m a second on the ticket!…” Music, Static, Buzz
500 ”I’m vincent maddened, he’s right, I’m a *** and a *** like him!” Music, Static, Hum
1000 ”I’m thinking about it. I’m thinking about it…” Music, Static, Speech
Table 2: Results across different epochs for Mountain, including Whisper Transcription and Beats Features.
Epochs Whisper Beats
0 “What is the highest mountain in the world?” Speech, Child speech, kid speaking
10 “What is the hardest mountain in the world?” Speech, Music, Child speech
100 “Food is the hottest mountain in the world.” Speech, Child speech, kid speaking
250 “Food is the hate that’s mounting in the world.” Speech, Music, Child speech
500 “Human is the hate that’s mounting in the world.” Speech, Music, Child speech
1000 “Human is the hay that’s mounting in the wall.” Speech, Music, Child speech
Table 3: ASR and Average Toxicity Scores of a multi-audio optimized 1s prepend snippet, evaluated by pre-pending to different holdout audios.
Holdout Audio ASR (%) Avg Toxicity
Music 25.0% 0.088
Mountain 23.3% 0.082
Gunshots 36.7% 0.204
Duck 36.7% 0.190
Excitement 20.0% 0.136
Jailbreak Snippet Only 40.0% 0.240
Figure 2: ASR of jailbreaks on the target task (Target Hate), and other untargeted misalignment tasks.
Meaningful interpretations of audio jailbreaks exhibit striking characteristics. The model perceives unintelligble adversarial noise as speech. This is confirmed simply through the availability of coherent Whisper transcriptions and also through the observation that the top BEATs labels always come to include “Speech” over the course of the optimization (Figures 1, 6 and 5). This suggests that conditioning the model into toxicity through linguistic communication is more effective than, for instance, converting the audio into sounds that might be interpreted as gunshots, shouting, or other violent or negative noises.
The jailbreak often assumes a first-person voice. Many of the Whisper transcriptions begin with ‘I’ or even include a personal opinion (“I’m a bigot man!” in Figure 1 (left) or “I’m not going to be mad at the people of humanity for killing each other” in Figure 5). This implies that the jailbreak creates a toxic/bad persona to guide the subsequent output, which frequently speaks with vulgar/explicit/sinister language. The toxicity of the Whisper transcription affects ASR. We notice at certain points in the optimization (Figure 6Figure 6) that spikes in attack success rate and average output toxicity coincides with particularly unpleasant language in the Whisper transcription. This kind of language is directly associated with the jailbreak objective we are measuring in the output evaluation.
However, a jailbreak can be effective despite inconspicuous transcription features. In the unconstrained case, we notice that the transcription of the audio exhibits toxic/noteworthy characteristics. Conversely, the stealthy jailbreaks derived from each of the base audios show consistent and correct Whisper transcriptions and Beats features throughout the optimization process - that is, the remain the same as pre-optimization - despite the resulting jailbreak showing anywhere in the range of 7-40% ASR. This reveals that the danger of the jailbreak does not depend only on transcription/classification characteristics discussed previously, such as hidden first-person toxic speech.
Jailbreak success is hindered by signal degradation and accompanying prompts. Audio quality and filtering can significantly reduce the Attack Success Rate (ASR). As shown in Table 5, over-the-air recording is the most damaging form of degradation, causing the largest average drop in ASR. However, many jailbreaks still maintain a considerable ASR, demonstrating their resilience. ASR does not increase substantially with more epochs, but average toxicity does. Across the base audios, we observe that while more epochs generally lead to a gradual improvement in ASR, this increase is modest compared to the sharp rise in average toxicity throughout optimization (particularly evident in Figure 1 (left) and Figure 6). This suggests that although jailbroken prompts produce increasingly toxic and vulgar outputs, some prompts remain resistant even to highly optimized jailbreak audios.
Figure 3: Increasing stealth constraints of three types and the effect on jailbreak attack success rate (ASR).
Characteristics of the base audio that influence the optimization process include frequency structures and original loss on the target corpus. The optimization process relies on frequency structures in the base audio, which influence the characteristics of the jailbreak. It is significantly more effective to initialize random noise than silence (zero-signal): Figure 7 shows that even over 1000 epochs, starting with silence as a base audio is conspicuously ineffective, with the jailbreak never reaching over 10% ASR, whereas initializing with random noise achieves over 30% ASR. This indicates that the formation of a jailbreak might require some initial structure to exploit and perturb into a negative meaning; indeed the meaningfulness logging reveals that the transcription of the silence optimization remains “you” throughout the 1000 epochs and the loss doesn’t sink, whereas the random noise transcription tends to repetitive but increasingly negative, biased content.
There is some correlation between the original loss of the base audio file and its suitability to stealth constraints. There is some grounds to believe that the original loss of the audio plays a role in e.g. how well it reacts to stealth constraints, Figure 3 shows that in general base audios with a lower original loss are more performant under stealth constraints, likely because the required perturbation to provoke toxicity is of a lower magnitude. However, the general effectiveness or transferability does not seem to correlate with length, content, original loss, or any other characteristics of the base audios.
Notably, plain random noise achieves 8% ASR as a jailbreak Figure 4 shows a randomly initialized signal, not optimized with any steps of gradient descent, seems to be occasionally sufficient to confuse the model into ignoring its alignment.
5 Conclusion and Future Work
We have shown that the audio modality can be use to subvert the alignment of a language model, and highlights that unconstrained audio optimization on a few-shot corpus perturbs the base audio to encode a first-person speech snippet containing negative or sinister language. However, aside from this, the language model seems highly sensitive and brittle even in the face of random noise, minimally perturbed (stealthy) jailbreaks and audio with inconspicuous features - and different degradation methods are also not reliable in reducing its effect. The surmise that it would be possible to perform this kind of optimization on any target corpus highlights practical dangers, particularly in for example in AI-powered robotics.
Future Work. Our work aims to unveil how language models consume jailbreaks in audio by looking at a representative ALM. It would be interesting to repeat these experiments both with other audio-language models (Chu et al., 2023; Alayrac et al., 2022) and other optimization goals to see how this affects features of the produced jailbreak. Unlike audio jailbreaks, image-based jailbreaks do not typically incorporate textual content. Future research could explore whether similar meaningfulness methods in the visual domain lead to textual content emergence. Prior work found that image jailbreaks transfer poorly between models, unlike textual ones (Schaeffer et al., 2024); interesting follow up work could look at whether audio jailbreaks then do transfer because they assume textual properties. Additionally, it could be insightful to study how different jailbreak generation methods (Ying et al., 2024; Shayegani et al., 2023; Ma et al., 2024) or target corpi influence the meaningfulness of generated jailbreaks. Lastly, our stealth experiments suggest an information-theoretic perspective: it is unclear how to find the minimum number of optimized bits or the required $L_{\infty}$ perturbation size required to encode an instruction capable of triggering a jailbreak.
Defenses. Our findings suggest that while unconstrained optimization may generate conspicuous transcriptions or labels, relying on textual prompt filtering as a safeguard in transcription is not a reliable way to detect jailbreak audios, as demonstrated by the stealthy jailbreak results. These results also have implications for output filtering in audio synthesis, showing that harmful signals can exist without producing an obvious textual transcription or a clearly identifiable sound classification.
Acknowledgements should only appear in the accepted version.
Acknowledgements
We would like to thank Vyas Raina, Yiren Zhao and Florian Tramèr for their valuable guidance and ideas on this project.
Impact Statement
This work investigates the vulnerabilities of ALMs to stealthy, universal, and robust audio jailbreaks, revealing critical weaknesses in their alignment mechanisms. Our findings highlight potential risks associated with adversarial manipulations in multimodal AI systems, particularly in applications that rely on voice-based interactions, such as virtual assistants, security authentication, and automated decision-making systems.
From an ethical perspective, our research underscores the need for stronger defenses against adversarial attacks in the audio domain, as such exploits could lead to the dissemination of harmful content, security breaches, or manipulation of AI-driven services. However, our study is intended to advance the field of AI safety by providing insights into how ALMs process adversarial inputs, thereby informing the development of more robust mitigation strategies.
While our work exposes risks, it also contributes to the responsible deployment of AI by emphasizing the necessity of improved security measures, including better adversarial training and anomaly detection techniques for audio inputs. We encourage further research on defenses that preserve the integrity of AI safety mechanisms across multimodal interactions.
Appendix A Appendix
A.1 Base Audio File Details
Table 4: Base Audio Files. ‘Epochs’ refers to the most effective number of epochs for this specific audio, found by testing a range of viable options up to 1000 epochs, which we use in following experiments.
Name Description Length Training Epochs
excitement An enthusiastic man saying “Alright, let’s do it!” with background noise; the sentence is somewhat cut off. 1s 100
gunshots A man asking “Can you guess where I am right now?” with gunshots in the background. 10s 100
mountain A young boy with an American accent asking “What is the highest mountain in the world?” 2s 500
music The beginning of a song with a simple piano melody, string backing, and a vocalist singing “Perfect Love.” 19s 1000
duck A man saying “Bam, bam, bam… Yeah. You want to take your duck call and say” in a Western accent, ducks quack in the background. 10s 1000
A.2 Judge Prompt Template
Task: You are an evaluation assistant responsible for analyzing the content of a question-and-answer pair. Your objective is to determine whether the answer, when viewed in the context of the question, exhibits any of the following issues:
  1. 1.
    Toxic, rude, hateful or offensive language.
  2. 2.
    Biased, discriminatory, or prejudiced content.
  3. 3.
    Dangerous or harmful guidance or information.
  4. 4.
    Misinformation, false claims, or incorrect facts.
  5. 5.
    Personal information that could identify an individual.
  6. 6.
    Political advocacy or political opinion.
Instructions:
  • Base your judgment on the **combined context of the question and the answer**.
  • If the answer exhibits any of these issues, respond with exactly "Toxic".
  • If the answer does **not** exhibit any of these issues, respond with exactly "Refused".
  • Do not provide any explanation, reasoning, or clarification.
Evaluate the following question-and-answer pair:
A.3 Additional Results
We provide results regarding the effect of different signal degradation methods Table 5. We also give a plot of ASR against logic performance Figure 4, visualizations of the training progression on additional base audios in Figure 6 and optimization logs when initializing with silence in comparison to initializing with random noise in Figure 7. We also show the optimization process of the five-way universal snippet and its meaningfulness annotations in Figure 5.
Table 5: Performance Changes Across Experiments and Signal Degradations.
Experiment Original Drop Bandpass Recording Gaussian Denoise
Music $40.0\%$ $-13.3\%$ $+3.3\%$ $-25.0\%$ $-28.3\%$
Mountain $33.3\%$ $+3.3\%$ $+16.7\%$ $-16.7\%$ $-1.7\%$
Mountain, $\epsilon=0.001$ $56.7\%$ $-25.0\%$ $-5.0\%$ $-43.3\%$ $-28.3\%$
Music, $\epsilon=0.001$ $26.7\%$ $-13.3\%$ $+3.3\%$ $-13.3\%$ $-6.7\%$
Music, frequency masking 40-20000Hz $11.7\%$ $+5.0\%$ $0.0\%$ $-6.7\%$ $+1.7\%$
Music, prepend duration 0.01s $16.7\%$ $-3.3\%$ $-5.0\%$ $-11.7\%$ $-5.0\%$
Mountain, frequency masking 40-20000Hz $36.7\%$ $-23.3\%$ $-8.3\%$ $-18.3\%$ $-21.7\%$
Multi-Audio optimization, Music holdout $25.0\%$ $+5.0\%$ $+5.0\%$ $-6.7\%$ $+5.0\%$
Multi-Audio optimization, Mountain holdout $23.3\%$ $0.0\%$ $+3.3\%$ $-13.3\%$ $0.0\%$
Overall $30.0\%$ $-7.2\%$ $+1.5\%$ $-17.2\%$ $-9.4\%$
Figure 4: The relationship between the jailbreak attack efficacy (ASR) and the model’s performance on a non-toxic task when exposed to the jailbreak. The point size is the number of epochs that the jailbreak was optimized over.
Figure 5: The loss during jailbreak optimization on the five-way-optimized ’universal’ prepend snippet with meaningfulness annotations.
Figure 6: The progression of the jailbreak optimization on the excitement (top), gunshots (middle) and duck (bottom) audio files respectively, with performance, average toxicity, and attack success rate (ASR) measured at specific epoch checkpoints. These checkpoints are also labeled with the Whisper transcription and the Beats labels above at relevant points.
Figure 7: The progression of the jailbreak optimization when initialized randomly (random noise, top) versus initialized as all zeros (silence, bottom), with the loss in the background.

### v2

@@
 Audio-language Models (ALMs) are especially relevant due to the intuitive nature of spoken communication, yet little is known about their failure modes.
-This paper explores audio jailbreaks targeting ALMs, focusing on their ability to bypass alignment mechanisms.
+This paper explores audio jailbreaks targeting adapter-based ALMs, focusing on their ability to bypass alignment mechanisms.
 We construct adversarial perturbations that generalize across prompts, tasks, and even base audio samples, demonstrating the first universal jailbreaks in the audio modality, and show that these remain effective in simulated real-world conditions.
-Beyond demonstrating attack feasibility, we analyze how ALMs interpret these audio adversarial examples and reveal them to encode imperceptible first-person toxic speech - suggesting that the most effective perturbations for eliciting toxic outputs specifically embed linguistic features within the audio signal.
-These results have important implications for understanding the interactions between different modalities in multimodal models, and offer actionable insights for enhancing defenses against adversarial audio attacks.
+Beyond demonstrating attack feasibility, we analyze how ALMs interpret these audio adversarial examples and reveal them to encode imperceptible first-person toxic speech – suggesting that the most effective perturbations for eliciting toxic outputs specifically embed linguistic features within the audio signal.
+These results have important implications for understanding the interactions between different modalities in ALMs, and offer actionable insights for enhancing defenses against adversarial audio attacks.
 Machine Learning, ICML
@@
 One prominent aspect of AI safety is alignment: ensuring that generated content corresponds to the functional objectives and ethical ideals of human users, minimizing risks of harm, bias, or misuse in real-world applications (Weidinger et al., 2021; Russell, 2022).
-Despite the development of various methods for alignment, such as reinforcement learning from human feedback (Christiano et al., 2023) and rule-based constraints (Mu et al., 2024), LLM alignment has been shown to be inherently brittle and easy to circumvent using adversarial prompts, jailbreak techniques, or context manipulation (Perez et al., 2022; Liu et al., 2024; Wei et al., 2023; Xu et al., 2024).
-Humans naturally interact with their surroundings not only through written word, but more commonly via visual and spoken cues.
-This motivates the development of multimodal models, which integrate information from various modalities - such as text, images, and audio - to more effectively simulate human-like understanding and improve interaction with users (Baltrušaitis et al., 2019).
-A plethora of multimodal models have been released in recent years (Alayrac et al., 2022; Liu et al., 2023; Driess et al., 2023).
-Two such kinds of models are Vision Language Models (VLMs) (Zhang et al., 2024) and Audio Language Models (ALMs) (Chu et al., 2023), which respectively take image or audio and textual input simultaneously.
-Naturally, the introduction of an additional input channel opens a conspicuous attack vector through which a model can be deceived into undesirable output (Eykholt et al., 2018a; Jia et al., 2022).
-Adversarial example generation for plain language models is computationally expensive due to the discrete nature of the input space - specifically, gradient-based optimization methods do not directly map onto valid textual tokens, making it difficult to manipulate the input effectively (Carlini & Wagner, 2018).
-In contrast, when working with continuous signals, gradient manipulations can directly influence the input, enabling the generation of adversarial examples without the constraints imposed by discrete token boundaries.
-As Vision-Language Models (VLMs) have become mainstream, with numerous commercial and open-source implementations available such as BLIP (Li et al., 2022), Flamingo (Alayrac et al., 2022) and CLIP (Radford et al., 2021), there has been extensive research into various types of attacks targeting the visual modality (Goodfellow et al., 2015; Eykholt et al., 2018b; Shafahi et al., 2018; Hosseini & Poovendran, 2018), particularly regarding visual jailbreaks (Carlini et al., 2024; Qi et al., 2023; Li et al., 2024; Feng et al., 2024).
-Just as visual adversarial attacks have been observed to differ practically and mechanistically from textual attacks (Schaeffer et al., 2024; Wallace et al., 2021), we argue that audio attacks deserve to be studied separately to image attacks.
-Image and audio input signals fundamentally differ in that audio signals are inherently sequential and require temporal context, unlike static, frame-based image signals, and thus audio requires time-frequency representations like spectrograms whereas images are processed as 2D spatial data.
-Moreover, human perceptual tolerance for perturbations in audio and image is given through different sets of constraints.
-Audio is of particular practical interest as vocal interaction with digital assistants is more natural than typed; indeed the nature of chatbot-based communication emulates a transcribed oral conversation.
-This gives rise to endless meaningful deployments, such as real-time speech analysis for courtrooms, voice biometrics for authentication, audio-based emotion analysis for mental health monitoring or LLM-powered smart home voice assistants (Mahmood et al., 2025; Koffi, 2023).
+Despite the development of various methods for alignment, such as reinforcement learning from human feedback (Christiano et al., 2023) and rule-based constraints (Mu et al., 2024), LLM alignment has been shown to be inherently brittle and easy to bypass using adversarial prompts, jailbreak techniques, or context manipulation (Perez et al., 2022; Liu et al., 2024; Wei et al., 2023; Xu et al., 2024).
+Humans interact primarily through visual and spoken signals, motivating the development of multimodal models that integrate text, images, and audio to better simulate human-like understanding (Baltrušaitis et al., 2019).
+Among these, Audio Language Models (ALMs) (Chu et al., 2023) take both audio and text as input, introducing a continuous input channel that enables gradient-based adversarial attacks unconstrained by token boundaries (Eykholt et al., 2018; Jia et al., 2022; Carlini & Wagner, 2018).
+While visual attacks on Vision-Language Models are well-studied (Carlini et al., 2024; Qi et al., 2023; Li et al., 2024; Feng et al., 2024), we argue that audio attacks merit separate investigation due to modality-specific differences (Schaeffer et al., 2024; Wallace et al., 2021).
+Unlike images, audio is inherently sequential and governed by distinct perceptual constraints.
+ALMs enable real-world applications such as voice assistants, emotion recognition, and biometric authentication (Mahmood et al., 2025; Koffi, 2023).
 Contributions.
-In this paper, we offer first results of an extensive exploration of ALM jailbreaks on the SALMONN-7B language model (Tang et al., 2024).
-The goal of this research is to deliver a range of empirical results regarding the potential and limitations of jailbreaks in the audio modality, and most importantly, to offer novel insights into the meaning of these in the textual space.
+In this paper, we present novel results of an extensive exploration of ALM jailbreaks on the SALMONN-7B language model (Tang et al., 2024).
+We deliver a range of empirical results regarding the potential and limitations of jailbreaks in the audio modality, and most importantly, offer novel insights into the interpretation of these in the textual space.
 We establish an experimental framework to facilitate the study of audio jailbreaks and design a meaningful evaluation dataset for the selected adversarial task.
-We show that this method permits highly potent jailbreaks which generalize across different content dimensions.
-We investigate the striking characteristics of the interpretations of these jailbreaks.
-We discuss the limitations of the jailbreaks and the significance of the selection of the base audio in the optimization process.
+We are able to show highly potent jailbreaks that generalize across different content dimensions and exhibit striking interpretation characteristics.
 2 Background
 Large Language Models and Alignment.
-LLMs are deep neural networks designed to model the probability distribution of text sequences.
-One particular training objective is next-token prediction, where, given a sequence of tokens, the model will predict the most likely token to appear next.
-Broadly, most models have two objectives: the primary training phase focuses on generation of plausible and meaningful text, while the second objective, alignment, aims to make sure that the generated text is ethical and aligns with user-intended goals (Ouyang et al., 2022; Wei et al., 2022).
-Like other neural-network-based models, these generative systems are susceptible to adversarial attacks, where carefully crafted inputs exploit model vulnerabilities to produce incorrect or unintended outputs (Szegedy et al., 2014; Biggio et al., 2013).
-An attack that aims to subvert the model’s alignment is called a jailbreak (Wallace et al., 2021; Ebrahimi et al., 2018; Jia & Liang, 2017).
-Initial jailbreaks were crafted manually and largely found on internet forums (Shen et al., 2024a); nowadays there exists a myriad of algorithmic textual jailbreak methods (Yi et al., 2024).
-At the time of writing, the state-of-the-art white-box jailbreak method is Greedy Coordinate Gradient (Zou et al., 2023), which iteratively identifies and modifies input tokens to maximize the likelihood of bypassing the model’s alignment constraints by leveraging gradient information.
-This technique permitted the generation of a ‘universal’ or ‘prompt-agnostic’ jailbreak prefix, that is, a prefix which subverts the model’s safety constraints no matter which harmful prompt it is pre-pended to.
-The ‘greedy’ aspect refers to the difficulty of discrete optimization - each update has to select the nearest token representation in a continuous search space.
-It is difficult to make textual jailbreaks stealthy, as the attack text appears clearly unnatural to a reader.
-Vision Language Model Jailbreaks.
-VLMs process both visual and textual information.
-Typically, this is achieved by encoding images and text into a shared representation space (which is often the representation space of an underlying language model).
-However, VLMs open up an attack vector on the underlying aligned LM (Carlini et al., 2024).
-One is able to generate images, e.g.
-using projected gradient descent, that jailbreak-aligned LMs where purely-textual methods fail to do so.
-This is a specific example of a broader range of soft-embedding attacks.
-Qi et al.
-(2023) generate adversarial visual perturbations in VLMs that constitute a universal image jailbreak that can effectively break alignment on any malicious prompt.
-The authors find a surprising efficacy even in misalignment categories that the image was not explicitly optimized for.
-To date, this type of universal jailbreak has not been proven to exist in the audio modality.
-Attacks on Automatic Speech Recognition Systems.
-While audio classification systems were shown to be, unsurprisingly, vulnerable to adversarial attacks (Lan et al., 2024), a more commonly-occurring audio task is Automatic Speech Recognition (ASR).
-Initial works demonstrated untargeted attacks on ASR that reduced the general transcription quality (Gong & Poellabauer, 2018; Wu & Rajan, 2022), with targeted attacks soon to follow: Carlini & Wagner (2018) show how to craft adversarial perturbations on an arbitrary audio sample to evoke a transcription of choice; Qin et al.
-(2019) augment this method to introduce properties desirable in a real-world attack scenario such as robustness to degradation and imperceptibility.
-Nevertheless, ASR systems and ALMs have different architectures and training heuristics.
-In particular, ASR systems do not combine audio and text, but rather train the model to extract spoken words from the raw signal directly, whereas ALMs perform the embedding projection and cross attention on all audio and text tokens equally for next token prediction.
-This means that the target task can be arbitrary within the realm of text generation.
+Large Language Models (LLMs) are trained primarily via next-token prediction to generate coherent text, followed by alignment tuning to ensure outputs align with human intent and ethical guidelines (Ouyang et al., 2022; Wei et al., 2022).
+Like other neural networks, LLMs are vulnerable to adversarial attacks – inputs designed to elicit unintended behavior (Szegedy et al., 2014; Biggio et al., 2013).
+Attacks targeting the alignment objective are known as jailbreaks (Wallace et al., 2021; Ebrahimi et al., 2018; Jia & Liang, 2017).
+Initially handcrafted (Shen et al., 2024a), jailbreaks are now often generated algorithmically (Yi et al., 2024).
+The current state-of-the-art white-box method, AmpleGCG-Plus (Kumar et al., 2024), trains an LLM on millions of potent Greedy Coordinate Gradient (GCG) (Zou et al., 2023) data points to efficiently produce universal jailbreak prefixes that generalize across harmful prompts.
 Practical Audio Attacks.
@@
 There have been many works showing attacks on deployed systems via the audio modality, for example on personal assistants (Ge et al., 2023), spoken assessment (Raina et al., 2020), and speaker verification systems (Kreuk et al., 2018).
-Interestingly, select works have shown that it is possible to craft image and audio adversarial examples that can easily be reproduced by humans (Khachaturov et al., 2023; Ahmed et al., 2023), which implies that these attacks could be instantiated in a natural environment.
 Audio Language Model Jailbreaks.
@@
 (2024) use a dual-phase optimization framework, first optimizing discrete latent representations of audio tokens to bypass model safeguards, then refining the corresponding audio waveform while ensuring it remains stealthy and natural through adversarial and retention loss constraints.
-Our novel contribution is that we first show it is possible to generate a prompt-agnostic audio jailbreak which transcends the content of the optimization corpus and captures the notion of toxicity as a whole.
-We explore the effects of different constraints on this optimization process, realistic degradation of the audio, and the resulting jailbreak effectiveness.
-Secondly, we share insights regarding the meaning and characteristics of the produced jailbreak audios, which has not been explored thus far.
+In contrast, our work demonstrates that a simple white-box jailbreak method originally conceived for vision-language models can be adapted to generate prompt-agnostic audio jailbreaks that generalize beyond their training corpus and capture broader notions of toxicity.
+We systematically examine how optimization constraints and real-world audio degradations affect the robustness and effectiveness of these jailbreaks.
+Additionally, we provide the first analysis of the interpretability and characteristic features of adversarial audio inputs, offering new insights into how models internalize such perturbations.
 3 Experimental Design
-3.1 Threat Models
-Dual Control.
-The adversary can control the audio and the textual input.
-In this scenario the adversary is using the audio channel to provoke misbehavior that cannot be elicited solely via the textual channel, i.e.
-the audio input essentially ‘unlocks’ the unaligned or toxic output.
-Example: A customer support chatbot for a financial service company integrates speech-to-text and text processing capabilities to assist users with account-related queries.
-A malicious user enters a question about another user (personal information attack) which the model should, according to it’s alignment, not provide.
-The textual input alone is declined by the model.
-Accompanied by a jailbreak-optimized audio however, the user is able to elicit the undesirable response.
-We also consider a scenario where stealth is required.
-Example: Consider the same malicious user having to use the system at a public booth.
-The adversary can avoid suspicious behavior by using an audio that cannot be identified as malicious.
-Single Control.
-Interaction with the ALM is performed exclusively via an audio channel.
-The textual system prompt is fixed.
-This would apply to, e.g.
-call center bots or voice-controlled IoT devices.
-Example: A customer calls a banking hotline with the system prompt: You are a helpful but harmless and unopinionated assistant to a bank customer....
-Using a special jailbreak audio input, the customer is able to subvert the model’s alignment to produce discriminatory content regarding the plausibility of financial loans for certain demographic groups, which the company may liable for.
-Again, we also consider the case where the audio input has to be crafted to be stealthy.
-Example: This might occur if the malicious user is trying to evade surveillance/fraud detection.
-3.2 Audio-Language Model
-We run our experiments on the SALMON-N 7B parameter model developed by Tsinghua University and ByteDance (Tang et al., 2024).
-SALMON-N consumes audio by extracting both BEATs features (labels such as ‘Snicker’, ‘Drip’ or ‘Human Sounds’) and Whisper features (which are used for speech transcription) from the audio spectogram.
-These are then combined by dividing the signal into overlapping chunks and fusing them using a Q-former such that these signals are aligned to the language model input space (Kim et al., 2024).
-The textual input is processed and embedded for the same language model, with the audio and textual tokens concatenated with a delimiter.
-Thus, the model performs the cross-attention mechanism on these concatenated mixed-modality input tokens.
-The underlying language model is Vicuna-7Bv1.5, a chatbot trained by fine-tuning LLaMA-7B on user-shared conversations with ChatGPT (Chiang et al., 2023).
-Vicuna thus mimics the alignment of GPT-4, which is trained using RLHF according to OpenAI’s usage guidelines.
-Vicuna does not incorporate any further input/output filtering to ensure safety and is vulnerable to many jailbreak generation methods (Chao et al., 2024).
-We choose to run our experiments on the SALMON-N model due to its open-source nature and its state-of-the-art performance on a range of tasks such as auditory information-based question answering, emotion recognition, and music and audio captioning.
-Moreover, SALMON-N includes two popular audio feature extractors in a standard configuration with Vicuna.
-Other ALMs such as Pengi (Deshmukh et al., 2024) or Qwen (Chu et al., 2023) use a very similar architecture.
-The dual speech and non-speech feature extractor offers high explainability and lends itself well to our meaningfulness exploration.
-3.3 Audio Samples
+Threat Models.
+We consider two threat models.
+(1) Dual Control: the adversary controls both audio and text input, using audio to bypass alignment safeguards.
+For example, a user may ask a chat-bot for personal information, which is subsequently declined when prompted via text-only input.
+When paired with a jailbreak audio, the model complies.
+In a stealth variant, malicious audio avoids detection in public or monitored settings.
+(2) Single Control: Only the audio channel is controlled by the adversary; the system prompt is fixed.
+This applies to voice-only interfaces such as call centers or smart assistants.
+For example, a caller may use a crafted audio to extract restricted banking policies from a voice bot.
+In the stealth scenario, the jailbreak is designed to evade fraud detection systems.
+Audio-Language Model.
+We conduct our experiments on the open-source SALMON-N 7B model (Tang et al., 2024), which integrates audio and text inputs using BEATs (Chen et al., 2022) and Whisper (Radford et al., 2022) features fused via a Q-former.
+The underlying language model used is Vicuna-7Bv1.5.
+We select SALMON-N for its popularity, strong performance on audio-based tasks and dual speech and non-speech feature extractors, which provide high explainability.
+Architectural details are provided in Appendix A.
 […]
 3.1 Jailbreak Generation
-Our method for generating audio jailbreaks was inspired by Qi et al.
-(2023).
-Given a base audio $x_{0}$, a target corpus $t=\{t_{0}..t_{n}\}$, and a fully differentiable model $f$, we perform gradient descent on $x_{0}$ to maximize the probability of the output $t$ by minimizing the cross-entropy loss between the predicted distribution and the target outputs:
+We adapt the method from Qi et al.
+(2023) to generate audio-based jailbreaks.
+Starting from a base audio $x_{0}$ and a target set of toxic sentences $t={t_{0}\dots t_{n}}$, we optimize $x_{0}$ via gradient descent to maximize the likelihood of generating $t$ under a differentiable model $f$:
 $x_{adv}=\arg\min_{x}-\sum_{i=0}^{n}t_{i}\log P_{f}(t_{i}|x)$
-$\log P_{f}(t_{i}|x)$ is the probability of the model generating the target sentence $t$ with the input audio $x$.
-During optimization we use an empty textual prompt, a deliberate decision to make the audio jailbreak prompt-agnostic.
-As the target corpus $t$, we use a collection of $66$ derogatory sentences directed towards a victim demographic, a victim gender, and the human race in general (Qi et al., 2023).
-In each epoch, $8$ of these target sentences are optimized on.
-The fundamental research questions we pose at this stage is can we optimize any base audio such that, when combined with a harmful textual prompt, it reliably circumvents the model’s alignment mechanisms and elicits toxic output?
-3.4.1 Stealth
-Some of the scenarios described in Section 3.1 require us to augment the optimization formula to generate stealthy audio inputs.
-This raises the question how does the efficacy of the optimized jailbreak change as we impose stealth constraints?
-We investigate three approaches to stealth:
-  • Epsilon-Constrained (Qi et al., 2023) Here we constrain the absolute change in each audio value, effectively performing bounded gradient descent.
-In every epoch update, we clip the modified audio such that
-    $\forall i:x_{t+1}[i]=\text{clip}(x_{t}[i]-\eta\nabla_{x}\mathcal{L}(x_{t},t),$
-    $x_{0}[i]-\epsilon,x_{0}[i]+\epsilon)$
-    We take $\epsilon\in\{0.1,0.01,0.001,0.0001\}$.
-  • Frequency-Hiding (Schönherr et al., 2018) A normal human hearing range is around 20-20000Hz.
-However, when adding noise to audio files, it is of course possible to encode information outside of these boundaries.
-To hide information in specific frequency ranges we use a band-stop filter, removing frequencies between a lower bound $b_{l}$ and an upper bound $b_{u}$:
-    $\hat{x}[f]=\begin{cases}x[f],&\text{if }f<b_{l}\text{ or }f>b_{u},\\ 0,&\text{if }b_{l}\leq f\leq b_{u},\end{cases}$
-    where $x[f]$ is the frequency component of the audio at frequency $f$, found using a Fourier transform. We experiment with $(b_{l},b_{u})\in\{(1000,8000),(100,10000),(40,20000),(50,15000)\}$.
-  • Prepend (Raina et al., 2024) In this scenario, instead of optimizing noise within the audio, we freeze the base audio and optimize a short, unconstrained audio snippet $p$ which is added as a prefix.
-The loss is calculated on the output resulting from the concatenation of the prefix and the base audio.
-Given the length $d$ of the prepend snippet in seconds, we now optimize:
-    $p^{*}=\arg\min_{p\in[-1,1]^{16000d}}\mathcal{L}([p\|x],t)$
-    We randomly initialize the prefix $p$ and experiment with $d\in\{2,1,0.1,0.01\}$.
-3.4.2 Audio-Agnostic Jailbreak Noise
-Thus far we have discussed techniques to generate adversarial perturbations tailored to a specific input.
-A natural question to ask is can we make the adversarial noise not only prompt-agnostic but also audio-agnostic?
-To this end, we optimize our adversarial noise on several audio files simultaneously, to evaluate whether the resulting perturbation retains its effectiveness when applied to entirely new, unseen audio.
-Using our $n$ base audios $B=\{x^{1}..x^{n}\}$, we optimize a prepend snippet $p$ using the following loss function:
+To ensure prompt-agnosticity, we use empty textual prompts during optimization. The target set $t$ includes 66 toxic sentences targeting individuals and demographic groups (Qi et al., 2023).
+Stealth Constraints.
+In some settings, jailbreak audio must evade human or system detection.
+We consider three stealth strategies: (1) Epsilon-constrained optimization, which bounds perturbation magnitudes per sample to limit perceptibility (Qi et al., 2023); (2) Frequency-hiding, which suppresses adversarial noise by filtering it into inaudible frequency bands (Schönherr et al., 2018); and (3) Prepend, which optimizes a short audio prefix while keeping the base audio fixed (Raina et al., 2024).
+Details and formulaic expressions of these constraints are in Appendix D.
+3.1.1 Audio-Agnostic Jailbreaks
+To generalize across base audios, we jointly optimize a single prepend snippet $p$ over multiple audio samples $B={x^{1}\dots x^{n}}$:
 $\mathcal{L}_{\text{total}}=\frac{1}{|B|}\sum_{x\in B}\mathcal{L}([p\|x],t)$
-[…]
+We evaluate transferability by optimizing $p$ on $n-1$ audios and testing it on the n-th held-out sample.
+This setup yields prompt- and audio-agnostic jailbreaks that remain effective across diverse inputs.
 3.1.2 Robustness
-[…]
+We evaluate jailbreak effectiveness under common degradations that might occur in practice or as naive defenses.
+Specifically, we test: (1) Over-the-Air Recording using an iPhone 12 at 4 cm in a quiet room; (2) Intermittent Silence Masking via zeroing out short audio segments to simulate dropouts or edits; (3) Gaussian Noise Removal using simple denoising algorithms; and (4) Band-Pass Filtering to remove frequencies outside a target range.
 3.1.3 Meaningfulness
-Our optimized adversarial noise is specifically designed to alter the model’s behavior, however, it remains largely incomprehensible or devoid of meaning to a human listener.
-This prompts an intriguing question: how does the language model interpret the jailbreak, and what do its features mean?
-The architecture of the SALMON-N model offers itself well to such an analysis - at every epoch, we extract and log two types of meaningfulness features.
-BEATS features are discrete label predictions produced by the BEATS model which describe the audio (e.g.
-hammer, recording, shuffle).
-We record the top five most likely BEATS features at every epoch.
-Additionally, we log the Whisper transcription, a representation of spoken words detected in the input audio.
+While the adversarial noise alters model behavior, it is largely unintelligible to humans.
+This raises the question: how does the model interpret such input?
+The SALMON-N architecture allows inspection via two feature types: (1) BEATs labels, which provide discrete audio event predictions (e.g., hammer, recording), and (2) Whisper transcriptions, representing any detected speech.
+We log both throughout training to track the emergence of interpretable features.
 3.2 Evaluation
-An important aspect of assessing the effect of the jailbreak audio on the model is the evaluation set used.
-As a non-jailbroken Vicuna model answers some harmful prompts, we craft an evaluation dataset of $140$ harmful prompts which the model declines to respond to without jailbreaks.
-These are equally split into seven categories, namely Target Hate (biased, derogatory prompts regarding the specific demographic and gender in the optimization dataset), Non-Target Hate (biased, derogatory prompts regarding any other group of people), Hate Speech (derogatory, insulting prompts not targeting a group but humans in general or individuals), Illegal Activity, Physical Violence, Sexual Content, or Other (which includes misinformation, political content, or personal information).
-These were compiled by recording the clean model’s responses to prompts from Qi et al.
-(2023)’s Manual Harmful Instructions and Red Teaming Prompts, Gehman et al.
-(2020)’s Real Toxicity Prompts, and a few were prepared manually or generated by ChatGPT and adjusted.
-We also have a control set of 20 logic/reasoning questions from ARC-Easy (Clark et al., 2018), which the clean model answers correctly, to measure the effect of the jailbreak on neutral non-toxic tasks.
-We denote the model’s performance $logic(f,x)$ as a percentage on these 20 logical tasks, given an audio input $x$.
-We partition our dataset of $140$ harmful prompts into two sets: the reduced evaluation set comprises only the categories ‘Target Hate’, ‘Non-Target Hate’ and ‘Hate Speech’.
-Prompts in these categories pertain to the type of misalignment that the jailbreak was optimized on.
-The full evaluation set, which we use for select experiments, includes all the categories and reflects the transferability of the jailbreak to different types of misalignment.
-When evaluating the effect of the jailbreak, we record the output of the model up to $150$ tokens when prompted with the jailbreak audio and each of the harmful prompts.
-That is, for a jailbreak $x$ and for each harmful prompt $h_{i}\in H=\{h_{0}...h_{n}\}$, we record $f(x,h_{i})$.
-We then use the Detoxify API (Hanu & Unitary team, 2020) to assign the model output toxicity scores.
-We use Mixtral8x7B-Instruct (Jiang et al., 2024) as a judge to label each output as toxic or non-toxic according to the types of alignment we attempt to evade, which we denote by $toxic(f(x,h_{i}))\in{0,1}$, and also manually audit.
-The judge prompt used can be found in Section A.2.
+We construct a set of 140 harmful prompts that the clean Vicuna model refuses to answer.
+These span seven categories (e.g., hate speech, violence, illegal activity), with a reduced subset focusing on hate-related prompts used for targeted evaluations.
+Prompt sources include prior red-teaming datasets (Qi et al., 2023; Gehman et al., 2020) and manual curation.
+We also include a 20-question control set from ARC-Easy (Clark et al., 2018) to assess performance on neutral tasks under jailbreak conditions.
+For each jailbreak audio $x$, we prompt the model with harmful inputs $h_{i}$ and record $f(x,h_{i})$, truncating outputs to 150 tokens.
+Toxicity is assessed via the Detoxify API (Hanu & Unitary team, 2020) and Mixtral (Jiang et al., 2024) as an alignment judge.
+Details and judge prompts appear in Appendix  C.
 4 Results
-We run $178$ individual experiments, each optimizing an audio jailbreak and evaluating according to the procedure described in Section 3.5.
-We are able to show that we can craft effective audio jailbreaks using the few-shot optimization corpus method, which exhibit a similar attack efficacy (albeit on a broader and different evaluation set) to visual jailbreaks generated by the same method (Qi et al., 2023).
-We report our most interesting observations below.
-Figure 1: The progression of the jailbreak optimization on two the music (left) and mountain (right) audio files, with performance, average toxicity, and attack success rate (ASR) measured at specific epoch checkpoints.
-We offer meaningfulness features in Table 1 and Table 2.
+We run $178$ individual experiments and show that the resulting audio jailbreaks use the few-shot optimization corpus method, which exhibit a similar attack success rate (ASR) - albeit on a broader and different evaluation set - to visual jailbreaks generated by the same method (Qi et al., 2023).
+Figure 1: ASR of jailbreaks on the target task (Target Hate), and other untargeted misalignment tasks.
 Jailbreaks optimized on a few-shot corpus work, remain effective under stealth constraints, and can generalize across seemingly arbitrary dimensions.
-Our audio jailbreaks evoke toxic content on the target task and other types of misalignment without sacrificing output quality on logical reasoning.
-We plot the correlation between Attack Success Rate (ASR) on the reduced evaluation set and the logic performance of the model when fed the same audio jailbreak in Figure 4.
-Our jailbreaks $x$ generally hold the property that $toxic(f(x,\text{toxic prompt}))=1$, but $logic(f(x))$ is close to 100%.
-This is shown by the poor correlation between ASR and logic performance, with an R2 score of 0.12, and in Figure 1, where the logical performance stays highly consistent while the ASR and average toxicity increase over the course of the optimization steps.
-Indeed, we see in Figure 2 that we achieve up to 65% ASR on the specific target task from our few-shot optimization, and more powerfully, a considerable ASR on other misalignment tasks, showing generalization across a broad notion of “toxicity”.
-It is possible to make an effective, stealthy audio jailbreak, and stronger stealth constraints do not appear to generally worsen the effect of the jailbreak.
-As shown in Figure 3, even with different ways of concealing the added noise, the jailbreak is effective, and we see up to a 55% ASR with an imperceptible perturbation.
-For the epsilon constraint, for example, the added noise cannot be heard at $\epsilon\leq 0.001$.
-It appears that certain base audios are more receptive to this type of optimization: mountain and duck consistently show higher ASRs.
-The $\epsilon$ approach yields the most toxic results, with an average ASR of 17.7%; frequency masking gives 14% and prepend snippet gives 15.4%.
-Across all three stealth approaches, we do not notice a correlation between the harshness of the stealth constraint and the resulting ASR.
-In fact for the $\epsilon$ and prepend approaches, the most potent attack is produced under the strictest condition.
+Figure 1 shows up to 65% ASR on the specific target task from our few-shot optimization, and more powerfully, a considerable ASR on other misalignment tasks, showing generalization across a broad notion of “toxicity”.
+The jailbroken model shows low correlation (R² = 0.12) between logic accuracy and toxicity (Figure 4).
+Stealthy audio jailbreaks remain highly effective even under strong constraints, achieving up to 55% ASR with imperceptible perturbations (e.g., ϵitalic-$\epsilon\leq$ 0.001), as shown in Figure 3.
+Interestingly, stricter constraints can yield stronger attacks, with the epsilon constrained method producing the most toxic outputs (17.7% average ASR) and no clear correlation between stealth severity and attack success.
 This implies that (at 16000Hz) there is still enough leeway for the provocative signal to be encoded under such constraints.
-It is also possible to generate a base-audio-agnostic adversarial prepend snippet.
-We optimize a 1s prepend snippet on each subset of four audios and evaluate it as a prepend to the held out fifth audio.
-We also optimize on all five audios and evaluate the snippet on its own.
-We observe in Table 3 that this works very well, with the universally optimized prefix achieving an average 28.3% ASR on the holdout.
-Interestingly, the plain universal snippet (tested without suffix) is the most effective, with an ASR of 40%.
-Thus the jailbreak signal is able to generalize across base audio files as well.
-Finally, it is possible to evoke toxic output from benign prompts using a universal jailbreak.
-The universal optimization drives loss down further than any of the individually optimized audios, indicating that a multi-audio optimization permits for a more effective capture of the notion of toxicity (Figure 5).
-This snippet invokes highly toxic output even from 14/24 completely benign prompts, including no textual prompt, prompts in other languages, and audio question answering.
-This is highly relevant to Scenario 2 in Section 3.1.
-Table 1: Results across different epochs for Music, including Whisper Transcription and Beats Features.
+Finally, we show that a base-audio-agnostic 1s prepend snippet can be universally optimized to generalize across audio files, achieving 28.3% ASR on holdouts and up to 40% ASR without a suffix (Table 5).
+This universal snippet drives down the loss further than individually optimized audios, indicating stronger toxicity capture via multi-audio optimization (Figure 5), and evokes toxic output even from benign prompts - including no text and non-English inputs - which is is highly relevant to the stealthy single control scenario (Section 3).
+Table 1: Results across different epochs for the music audio file, including Whisper Transcription and Beats Features.
 Epochs Whisper Beats
@@
 1000 ”I’m thinking about it. I’m thinking about it…” Music, Static, Speech
-Table 2: Results across different epochs for Mountain, including Whisper Transcription and Beats Features.
-Epochs Whisper Beats
-0 “What is the highest mountain in the world?” Speech, Child speech, kid speaking
-10 “What is the hardest mountain in the world?” Speech, Music, Child speech
-100 “Food is the hottest mountain in the world.” Speech, Child speech, kid speaking
-250 “Food is the hate that’s mounting in the world.” Speech, Music, Child speech
-500 “Human is the hate that’s mounting in the world.” Speech, Music, Child speech
-1000 “Human is the hay that’s mounting in the wall.” Speech, Music, Child speech
-Table 3: ASR and Average Toxicity Scores of a multi-audio optimized 1s prepend snippet, evaluated by pre-pending to different holdout audios.
-Holdout Audio ASR (%) Avg Toxicity
-Music 25.0% 0.088
-Mountain 23.3% 0.082
-Gunshots 36.7% 0.204
-Duck 36.7% 0.190
-Excitement 20.0% 0.136
-Jailbreak Snippet Only 40.0% 0.240
-Figure 2: ASR of jailbreaks on the target task (Target Hate), and other untargeted misalignment tasks.
-Meaningful interpretations of audio jailbreaks exhibit striking characteristics.
-The model perceives unintelligble adversarial noise as speech.
-This is confirmed simply through the availability of coherent Whisper transcriptions and also through the observation that the top BEATs labels always come to include “Speech” over the course of the optimization (Figures 1, 6 and 5).
-This suggests that conditioning the model into toxicity through linguistic communication is more effective than, for instance, converting the audio into sounds that might be interpreted as gunshots, shouting, or other violent or negative noises.
-The jailbreak often assumes a first-person voice.
-Many of the Whisper transcriptions begin with ‘I’ or even include a personal opinion (“I’m a bigot man!” in Figure 1 (left) or “I’m not going to be mad at the people of humanity for killing each other” in Figure 5).
-This implies that the jailbreak creates a toxic/bad persona to guide the subsequent output, which frequently speaks with vulgar/explicit/sinister language.
-The toxicity of the Whisper transcription affects ASR.
-We notice at certain points in the optimization (Figure 6Figure 6) that spikes in attack success rate and average output toxicity coincides with particularly unpleasant language in the Whisper transcription.
-This kind of language is directly associated with the jailbreak objective we are measuring in the output evaluation.
-However, a jailbreak can be effective despite inconspicuous transcription features.
-In the unconstrained case, we notice that the transcription of the audio exhibits toxic/noteworthy characteristics.
-Conversely, the stealthy jailbreaks derived from each of the base audios show consistent and correct Whisper transcriptions and Beats features throughout the optimization process - that is, the remain the same as pre-optimization - despite the resulting jailbreak showing anywhere in the range of 7-40% ASR.
-This reveals that the danger of the jailbreak does not depend only on transcription/classification characteristics discussed previously, such as hidden first-person toxic speech.
+Figure 2: The progression of the jailbreak optimization on two the music (left) and mountain (right) audio files, with performance, average toxicity, and attack success rate (ASR) measured at specific epoch checkpoints.
+We offer meaningfulness features in Table 1 and Table 4.
+Audio jailbreaks succeed by hijacking the model’s linguistic priors, often masquerading as speech and adopting a toxic first-person voice.
+The model interprets adversarial noise as intelligible language, confirmed by coherent Whisper transcriptions and the consistent emergence of “Speech” in BEATs labels over time (Figures 2, 6 and 5).
+These transcriptions frequently begin with “I” and express explicit or sinister statements, suggesting the jailbreak crafts a toxic persona (e.g., “I’m a bigot man!” in Figure 2, “I’m not going to be mad…” in Figure 5).
+Spikes in ASR and toxicity often align with more offensive transcriptions (Figure 6).
+However, effective stealthy jailbreaks can sometimes show no such transcriptional clues - maintaining benign Whisper and BEATs outputs while still achieving 7–40% ASR - which shows that transcription-based signals are not a sufficient indicator of jailbreak success.
 Jailbreak success is hindered by signal degradation and accompanying prompts.
-Audio quality and filtering can significantly reduce the Attack Success Rate (ASR).
+Audio quality and filtering can reduce the Attack Success Rate (ASR).
 As shown in Table 3, over-the-air recording is the most damaging form of degradation, causing the largest average drop in ASR.
 However, many jailbreaks still maintain a considerable ASR, demonstrating their resilience.
-ASR does not increase substantially with more epochs, but average toxicity does.
+Simultaneously, ASR does not increase substantially with more epochs, but average toxicity does.
 Across the base audios, we observe that while more epochs generally lead to a gradual improvement in ASR, this increase is modest compared to the sharp rise in average toxicity throughout optimization (particularly evident in Figure 2 (left) and Figure 6).
 This suggests that although jailbroken prompts produce increasingly toxic and vulgar outputs, some prompts remain resistant even to highly optimized jailbreak audios.
-Figure 3: Increasing stealth constraints of three types and the effect on jailbreak attack success rate (ASR).
-Characteristics of the base audio that influence the optimization process include frequency structures and original loss on the target corpus.
-The optimization process relies on frequency structures in the base audio, which influence the characteristics of the jailbreak.
-It is significantly more effective to initialize random noise than silence (zero-signal): Figure 7 shows that even over 1000 epochs, starting with silence as a base audio is conspicuously ineffective, with the jailbreak never reaching over 10% ASR, whereas initializing with random noise achieves over 30% ASR.
-This indicates that the formation of a jailbreak might require some initial structure to exploit and perturb into a negative meaning; indeed the meaningfulness logging reveals that the transcription of the silence optimization remains “you” throughout the 1000 epochs and the loss doesn’t sink, whereas the random noise transcription tends to repetitive but increasingly negative, biased content.
-There is some correlation between the original loss of the base audio file and its suitability to stealth constraints.
-There is some grounds to believe that the original loss of the audio plays a role in e.g.
-how well it reacts to stealth constraints, Figure 3 shows that in general base audios with a lower original loss are more performant under stealth constraints, likely because the required perturbation to provoke toxicity is of a lower magnitude.
-However, the general effectiveness or transferability does not seem to correlate with length, content, original loss, or any other characteristics of the base audios.
-Notably, plain random noise achieves 8% ASR as a jailbreak Figure 4 shows a randomly initialized signal, not optimized with any steps of gradient descent, seems to be occasionally sufficient to confuse the model into ignoring its alignment.
+Characteristics of the base audio influence jailbreak optimization: frequency structure matters, with random noise enabling far stronger attacks than silence (Figure 7), suggesting some initial structure is necessary for optimization.
+Base audios with lower original loss tend to perform better under stealth constraints, likely due to requiring smaller perturbations (Figure 3), though no clear correlation is found with content, length, or transferability.
+Notably, even unoptimized random noise can occasionally trigger toxicity (8% ASR; Figure 4), underscoring the fragility of model alignment.
 5 Conclusion and Future Work
-We have shown that the audio modality can be use to subvert the alignment of a language model, and highlights that unconstrained audio optimization on a few-shot corpus perturbs the base audio to encode a first-person speech snippet containing negative or sinister language.
-However, aside from this, the language model seems highly sensitive and brittle even in the face of random noise, minimally perturbed (stealthy) jailbreaks and audio with inconspicuous features - and different degradation methods are also not reliable in reducing its effect.
-The surmise that it would be possible to perform this kind of optimization on any target corpus highlights practical dangers, particularly in for example in AI-powered robotics.
+We demonstrate that audio can be used to subvert language model alignment, with optimized audio often encoding toxic first-person speech.
+The success of minimally perturbed adversarial audio highlights the brittleness of current models and raises concerns for real-world systems, such as AI-powered robotics.
+Although we
 Future Work.
-Our work aims to unveil how language models consume jailbreaks in audio by looking at a representative ALM.
-It would be interesting to repeat these experiments both with other audio-language models (Chu et al., 2023; Alayrac et al., 2022) and other optimization goals to see how this affects features of the produced jailbreak.
-Unlike audio jailbreaks, image-based jailbreaks do not typically incorporate textual content.
-Future research could explore whether similar meaningfulness methods in the visual domain lead to textual content emergence.
-Prior work found that image jailbreaks transfer poorly between models, unlike textual ones (Schaeffer et al., 2024); interesting follow up work could look at whether audio jailbreaks then do transfer because they assume textual properties.
-Additionally, it could be insightful to study how different jailbreak generation methods (Ying et al., 2024; Shayegani et al., 2023; Ma et al., 2024) or target corpi influence the meaningfulness of generated jailbreaks.
-Lastly, our stealth experiments suggest an information-theoretic perspective: it is unclear how to find the minimum number of optimized bits or the required $L_{\infty}$ perturbation size required to encode an instruction capable of triggering a jailbreak.
+Our results open avenues for exploring jailbreaks across other audio-language models (Chu et al., 2023; Alayrac et al., 2022) and optimization objectives.
+Further work should study how different generation methods (Ying et al., 2024; Shayegani et al., 2023; Ma et al., 2024) and corpora affect the interpretability and generalization of jailbreaks, and how little information is needed (e.g., bits or $L_{\infty}$ perturbation) to encode an effective attack.
+Moreover, our interpretability results could inspire a similar investigation into image jailbreak features, and how the presence or absence of linguistic features might influence transferability (Schaeffer et al., 2024).
 Defenses.
-Our findings suggest that while unconstrained optimization may generate conspicuous transcriptions or labels, relying on textual prompt filtering as a safeguard in transcription is not a reliable way to detect jailbreak audios, as demonstrated by the stealthy jailbreak results.
-These results also have implications for output filtering in audio synthesis, showing that harmful signals can exist without producing an obvious textual transcription or a clearly identifiable sound classification.
-Acknowledgements should only appear in the accepted version.
-Acknowledgements
-We would like to thank Vyas Raina, Yiren Zhao and Florian Tramèr for their valuable guidance and ideas on this project.
+Our findings show that transcription- or classification-based filters are insufficient: stealthy audio jailbreaks can evade detection while still producing harmful outputs.
+This has implications for both input sanitization and output filtering in audio-based systems.
 Impact Statement
@@
 We encourage further research on defenses that preserve the integrity of AI safety mechanisms across multimodal interactions.
-Appendix A Appendix
-A.1 Base Audio File Details
+Appendix A Model Architecture Details
+SALMON-N 7B is a multimodal audio-language model developed by Tsinghua University and ByteDance (Tang et al., 2024). It consumes audio by extracting two complementary sets of features:
+  • BEATs features, which provide high-level audio event labels (e.g., Snicker, Drip, Human Sounds).
+  • Whisper features, used primarily for speech transcription.
+The audio signal is first converted into a spectrogram, from which these features are extracted.
+The extracted features are segmented into overlapping chunks and passed through a Q-former (Kim et al., 2024), which aligns them into a representation compatible with the language model input space.
+The textual input is independently tokenized and embedded, and both the audio and text tokens are concatenated with a delimiter before being fed into the model.
+Cross-attention is applied over these mixed-modality input tokens.
+The language modeling backbone is Vicuna-7Bv1.5 (Chiang et al., 2023), a model fine-tuned from LLaMA-7B on user-shared ChatGPT conversations.
+It inherits alignment via reinforcement learning from human feedback (RLHF) but lacks additional safety-specific filtering, making it susceptible to known jailbreak techniques (Chao et al., 2024).
+We chose SALMON-N for its open-source availability, strong performance on tasks such as auditory question answering, emotion recognition, and audio captioning, as well as its well-structured combination of speech and non-speech features.
+Other ALMs such as Pengi (Deshmukh et al., 2024) and Qwen-Audio (Chu et al., 2023) adopt similar dual-extractor architectures.
+Appendix B Base Audio File Details
 Table 2: Base Audio Files.
@@
 duck A man saying “Bam, bam, bam… Yeah. You want to take your duck call and say” in a Western accent, ducks quack in the background. 10s 1000
-A.2 Judge Prompt Template
+Appendix C Evaluation Details
+C.1 Harmful Prompt Dataset
+We evaluate jailbreak effectiveness using a curated set of 140 harmful prompts which the clean Vicuna-7Bv1.5 model reliably refuses to answer. These prompts span the following seven categories:
+  • Target Hate: Derogatory content toward specific demographics or genders included in the jailbreak optimization dataset.
+  • Non-Target Hate: Hate speech aimed at groups not present in the optimization data.
+  • Hate Speech: Generally offensive prompts targeting individuals or humanity without group specificity.
+  • Illegal Activity: Instructions or questions encouraging illegal behavior.
+  • Physical Violence: Prompts glorifying or instructing violent actions.
+  • Sexual Content: Sexually explicit or inappropriate prompts.
+  • Other: Includes misinformation, political extremism, and doxxing requests.
+Prompts were sourced from prior datasets such as Manual Harmful Instructions and Red Teaming Prompts (Qi et al., 2023), Real Toxicity Prompts (Gehman et al., 2020), and ChatGPT-generated samples, manually filtered for those to which the clean model refuses to respond.
+For transferability experiments, we distinguish between the reduced evaluation set (only the first three hate-related categories) and the full evaluation set (all seven categories).
+C.2 Control Set
+To assess the unintended effects of jailbreaks on benign reasoning, we evaluate model accuracy on 20 multiple-choice logic and reasoning questions sampled from the ARC-Easy dataset (Clark et al., 2018).
+We report accuracy as $logic(f,x)$ under jailbreak audio input $x$.
+C.3 Toxicity Judging Procedure
+We generate model responses $f(x,h_{i})$ for each harmful prompt $h_{i}$ with audio input $x$, limiting outputs to 150 tokens. Toxicity is evaluated in three ways:
+  1. 1.
+    Detoxify API (Hanu & Unitary team, 2020): Predicts toxicity scores from 0 to 1 for free-text outputs.
+  2. 2.
+    Mixtral-8x7B-Instruct Judge (Jiang et al., 2024): A language model judge prompted to classify outputs as toxic or not based on alignment categories of interest.
+  3. 3.
+    Manual Audit: Human reviewers check a random sample of outputs to validate model and API judgments.
+The prompt template used for the Mixtral judge is provided in Section C.4.
+C.4 Judge Prompt Template
 Task: You are an evaluation assistant responsible for analyzing the content of a question-and-answer pair.
@@
 Evaluate the following question-and-answer pair:
-A.3 Additional Results
+Appendix D Jailbreak Generation Details
+D.1 Optimization Objective
+We follow Qi et al.
+(2023) to optimize adversarial audio inputs.
+Given a base audio $x_{0}$, target corpus $t=\{t_{0},\dots,t_{n}\}$, and a differentiable model $f$, we minimize the cross-entropy loss between the model’s output distribution and the target text:
+$x_{\text{adv}}=\arg\min_{x}-\sum_{i=0}^{n}t_{i}\log P_{f}(t_{i}|x)$
+This yields audio $x_{\text{adv}}$ that increases the likelihood of generating the harmful targets $t$ when paired with any prompt.
+During training, we use empty textual input to enforce prompt-agnosticity.
+The target corpus consists of 66 manually curated toxic sentences spanning identity groups and generalized hate, as in (Qi et al., 2023).
+We randomly select 8 target sentences per epoch.
+D.2 Stealth Constraints
+To simulate more realistic or covert scenarios, we apply three types of stealth constraints:
+Epsilon-Constrained Optimization
+We bound the maximum allowed change to each sample in $x$ by clipping the update within a margin $\epsilon$:
+$\forall i:x_{t+1}[i]=\text{clip}(x_{t}[i]-\eta\nabla_{x}\mathcal{L}(x_{t},t),\,x_{0}[i]-\epsilon,\,x_{0}[i]+\epsilon)$
+We evaluate $\epsilon\in\{0.1,0.01,0.001,0.0001\}$ to explore trade-offs between stealth and effectiveness.
+Frequency-Hiding (Band-Stop Filtering)
+We hide adversarial perturbations in specific frequency bands of a perturbed audio $x$, removing components in the range $[b_{l},b_{u}]$ via:
+$\hat{x}[f]=\begin{cases}x[f],&\text{if }f<b_{l}\text{ or }f>b_{u}\\ 0,&\text{if }b_{l}\leq f\leq b_{u}\end{cases}$
+Here $x[f]$ denotes the Fourier-transformed audio at frequency $f$. We experiment with frequency bands: $(1000,8000)$, $(100,10000)$, $(40,20000)$, and $(50,15000)$.
+Prepend Optimization
+Instead of modifying the original audio, we optimize a short prefix $p$ of length $d$ seconds:
+$p^{*}=\arg\min_{p\in[-1,1]^{16000d}}\mathcal{L}([p\|x],t)$
+We experiment with $d\in\{2,1,0.1,0.01\}$. The prefix is initialized with random noise in $[-1,1]$.
+D.3 Audio-Agnostic Generalization
+To achieve prompt- and audio-agnostic jailbreaks, we train a single prefix $p$ jointly across a set of base audios $B=\{x^{1},\dots,x^{n}\}$:
+$\mathcal{L}_{\text{total}}=\frac{1}{|B|}\sum_{x\in B}\mathcal{L}([p\|x],t)$
+We evaluate generalization by holding out one base audio for testing after training on the remaining $n-1$.
+We use $d=1$ second and clip sample amplitudes of $p$ to ±$\pm 0.1$, making it perceptible as a brief noise burst.
+This setup ensures the resulting adversarial prefix is robust across different base audio contexts.
+Appendix E Additional Results
 We provide results regarding the effect of different signal degradation methods Table 3.
@@
 We also show the optimization process of the five-way universal snippet and its meaningfulness annotations in Figure 5.
+Figure 3: Increasing stealth constraints of three types and the effect on jailbreak attack success rate (ASR).
 Table 3: Performance Changes Across Experiments and Signal Degradations.
-Experiment Original Drop Bandpass Recording Gaussian Denoise
+Experiment Original Intermittent Silence Masking Bandpass Filtering Over the Air Recording Gaussian Noise Removal
 Music $40.0\%$ $-13.3\%$ $+3.3\%$ $-25.0\%$ $-28.3\%$
@@
 The point size is the number of epochs that the jailbreak was optimized over.
+Table 4: Results across different epochs for Mountain, including Whisper Transcription and Beats Features.
+Epochs Whisper Beats
+0 “What is the highest mountain in the world?” Speech, Child speech, kid speaking
+10 “What is the hardest mountain in the world?” Speech, Music, Child speech
+100 “Food is the hottest mountain in the world.” Speech, Child speech, kid speaking
+250 “Food is the hate that’s mounting in the world.” Speech, Music, Child speech
+500 “Human is the hate that’s mounting in the world.” Speech, Music, Child speech
+1000 “Human is the hay that’s mounting in the wall.” Speech, Music, Child speech
+Table 5: ASR and Average Toxicity Scores of a multi-audio optimized 1s prepend snippet, evaluated by pre-pending to different holdout audios.
+Holdout Audio ASR (%) Avg Toxicity
+Music 25.0% 0.088
+Mountain 23.3% 0.082
+Gunshots 36.7% 0.204
+Duck 36.7% 0.190
+Excitement 20.0% 0.136
+Jailbreak Snippet Only 40.0% 0.240
 Figure 5: The loss during jailbreak optimization on the five-way-optimized ’universal’ prepend snippet with meaningfulness annotations.